# Tree-Like Shortcuttings of Trees

### v1

@@ 2 Preliminaries
Given a graph $G$, let $V(G)$ and $E(G)$ be the vertex and edge set of $G$. A tree decomposition of $G$ is a pair $(X,T)$ where $T$ is a tree and $X=\{X_{1},X_{2},\ldots,X_{l}\}$ is a set of subsets of $V(G)$, called bags, associated with nodes in $T$ such that:
  • Every vertex in $V(G)$ is in at least one bag in $X$.
  • For every edge $e$ in $E(G)$, there is at least one bag containing both endpoints of $e$.
  • For every vertex $u$ in $V(G)$, the bags containing $u$ in $X$ inducing a subtree of $T$.
The width of $(X,T)$ is $\max_{X_{i}\in X}|X_{i}|-1$. The treewidth of $G$ is the minimum width over all the tree decomposition of $G$.
The treewidth of a graph is closely related to the size of the complete graph minors it contains. Recall that a minor of a graph $G$ is any graph obtained from $G$ by a sequence of vertex deletions, edge deletions, and edge contractions. Let $K_{h}$ denote the complete graph on $h$ vertices. It is widely known that treewidth is monotone under taking minors. Then, we obtain the following fundamental fact:
Fact 2.1.
If a graph $G$ contains $K_{h}$ as a minor, then the treewidth of $G$ is at least h1h-1.
@@
The arboricity of a graph $G$ measures how sparse its subgraphs can be. Formally, the arboricity of $G$, denoted by $\mathrm{arb}(G)$, is defined as
$\mathrm{arb}(G)\;=\;\max_{H\subseteq G,\,|V(H)|\geq 2}\left\lceil\frac{|E(H)|}{|V(H)|-1}\right\rceil,$
where the maximum is taken over all subgraphs $H$ of $G$ with at least two vertices. Nash-Williams proved that the arboricity equal to the minimum number of forests that $G$ can decomposed into [NW64].
Spanner.
A $t$-spanner of a graph $G$ is a spanning subgraph $H$ that approximately preserves distances between all pairs of vertices. Formally, for every $u,v\in V(G)$, $d_{H}(u,v)\leq t\cdot d_{G}(u,v)$. A $t$-spanner of a metric space $(X,\delta)$ is a $t$ spanner of $\left(X,\binom{X}{2},\delta\right)$. A spanner of a metric space is often known as geometric spanner.
Tree cover.
A tree cover of a graph $G$ with stretch $t$ is a collection of spanning trees $\mathcal{T}$ of $G$ such that, for every $u,v\in V(G)$:
  • The distance between $u$ and $v$ in any tree $T\in\mathcal{T}$ is at least the distance in $G$, i.e., $d_{T}(u,v)\geq d_{G}(u,v)$.
  • There exists a tree $T\in\mathcal{T}$ such that $d_{T}(u,v)\leq t\cdot d_{G}(u,v)$.
Similar to geometric spanner, a tree cover of a metric space $(X,\delta)$ is a tree cover of the complete graph $\left(X,\binom{X}{2},\delta\right)$.
Compact routing scheme.
A routing scheme is a distributed algorithm that, given a packet with a designated source and destination, specifies how the packet is forwarded along the edges of the network. Each node in the network has its own routing table, which stores local routing information, as well as a unique label. During a preprocessing phase, the network is initialized so that each node is assigned a routing table and a label. In the labeled model, labels are chosen by the designer (typically of size $\mathrm{poly}(\log n)$), while in the name-independent model, labels are chosen adversarially.
@@
Following standard notions (e.g., [NS07, AS24]), we introduce the follwing Ackermann function.
Definition 2.1 (Ackermann functions).
For all $k\geq 0$, the functions $A(k,n)$ and $B(k,n)$ are defined as follows:
@@
Then, we have the following definition of the inverse Ackermann function:
Definition 2.2 (Inverse Ackermann function).
For all $k\geq 0$, the function $\alpha_{k}(n)$ is defined as follows:
$\alpha_{2k}(n):=\min\{\,s\geq 0:A(k,s)\geq n\,\},\quad\text{for all }n\geq 0,$
$\alpha_{2k+1}(n):=\min\{\,s\geq 0:B(k,s)\geq n\,\},\quad\text{for all }n\geq 0.$
@@
3 Bounded treewidth tree covers with small hop-diameter
In this section, we show tight tradeoff between treewidth and hop-diameter for 1-spanners of tree metrics. In particular, the upper bound (Section˜1.1.1) is proved in Section˜3.1 and the matching lower bound (Section˜1.1.1) is proved in Section˜3.2.
@@ 3.1 Upper bound
In this section we prove Section˜1.1.1.
@@ 3.1.1 Hop-diameter 2
For every tree metric $M_{T}=(T,\delta_{T})$ induced by a tree $T$ there is a $1$-spanner $H_{2}$ with hop-diameter 2 and treewidth $O(\log{n})$.
@@
Let $\mathcal{T}_{1},\ldots,\mathcal{T}_{g}$ be the tree decompositions of spanners constructed for the subtrees in $T\setminus v$. Create a new bag $B$ containing only the vertex $v$ and add $v$ to every bag in each $\mathcal{T}_{1},\ldots,\mathcal{T}_{g}$. The tree decomposition of $H_{2}$ is obtained by connecting $B$ to the roots of each of $\mathcal{T}_{1},\ldots,\mathcal{T}_{g}$. The treewidth satisfies recurrence $W_{2}(1)=0$ and $W_{2}(n)=W_{2}(n/2)+1$, which has solution $W_{2}(n)=O(\log{n})$.
@@ 3.1.2 Hop-diameter 3
For every tree metric $M_{T}=(T,\delta_{T})$ induced by a tree $T$ there is a $1$-spanner $H_{3}$ with hop-diameter 3 and treewidth $O(\log{n}/\log\log{n})$.
@@
Let $\mathcal{T}_{1},\ldots,\mathcal{T}_{g}$ be the tree decompositions of spanners constructed for the trees in $T\setminus X$. For every $\mathcal{T}_{i}$, let $u$ and $v$ be the two vertices from $X$ adjacent to the corresponding subtree $W_{i}$ in $T\setminus X$. (Lemma˜3.1 guarantees that there is at most two such vertices; it is possible that $u=v$.) Add $u$ and $v$ to each bag in $\mathcal{T}_{i}$. Construct a new bag $B$ containing all the vertices in $X$ and connect it to the roots of each $\mathcal{T}_{1},\ldots,\mathcal{T}_{g}$. The treewidth of $H_{3}$ satisfies $W_{3}(n)\leq n$ for $n\leq\ell_{3}$ and $W_{3}(n)\leq W_{3}(n/\ell_{3})+2$ for $n>\ell_{3}$. Recall that $\ell_{3}$ is fixed and does not change across different levels of recursion. The recurrence satisfies $W_{3}(n)=O(\log{n}/\log\log{n})$.
@@ 3.1.3 Hop-diameter $k\geq 4$
For every tree metric $M_{T}=(T,\delta_{T})$ induced by a tree $T$ and every $k\geq 4$ there is a $1$-spanner $H_{k}$ with hop-diameter $k$ and treewidth $O(k\log^{2/k}{n})$ for even values of $k$ and $O(k(\frac{\log{n}}{\log\log{n}})^{2/(k-1)})$ for odd values of $k$.
@@
Next, we analyze the treewidth of $H_{k}$. Let $\mathcal{T}_{1},\ldots,\mathcal{T}_{g}$ be the tree decompositions of spanners constructed for the trees in $T\setminus X$. For every $\mathcal{T}_{i}$, let $u$ and $v$ be the two vertices from $X$ adjacent to the corresponding subtree $W_{i}$ in $T\setminus X$. (Lemma˜3.1 guarantees that there is at most two such vertices; it is possible that $u=v$.) Add $u$ and $v$ to each bag in $\mathcal{T}_{i}$. Let $\mathcal{T}_{X}$ be the tree decomposition of $T_{X}$, where $T_{X}$ is defined as in the previous paragraph. Since $\mathcal{T}_{X}$ is a valid tree decomposition and $T_{X}$ contains an edge $(u,v)$, then $\mathcal{T}_{X}$ contains a bag $L_{u,v}$ with both $u$ and $v$. Connect the root of $\mathcal{T}^{\prime}$ to $L_{u,v}$. This concludes the description of the tree decomposition $\mathcal{T}$ of $T$. The treewidth of $\mathcal{T}$ satisfies the recurrence $W_{k}(n)=W_{k-2}(n)$ for $n\leq\ell_{k}$ and $W_{k}(n)\leq\max(W_{k-2}(\ell_{k}),W_{k}(n/\ell_{k})+2)$ otherwise.
@@
To argue that $H_{k}$ is a 1-spanner with hop-diameter $k$, consider the last step of the construction where both $u$ and $v$ were in the same tree $T$. If $u$ and $v$ were considered in the same base case, then $T$ is equipped with a recursive construction with parameter k2k-2. By induction, there is a $1$-spanner path between them with hop-diameter k2k-2. Otherwise, let $W_{u}$ (resp., $W_{v}$) be the subtree in $T\setminus X$ and let $u^{\prime}$ (resp., $v^{\prime}$) be the vertex in $X$ that is incident on $W_{u}$ (resp., $W_{v}$). By construction, there is a 1-spanner path between $u^{\prime}$ and $v^{\prime}$ with at most k2k-2 hops. (The cases where $u^{\prime}=u$ or $v^{\prime}=v$ are handled similarly.) ∎
@@ 3.2 Lower bound
We show the treewidth lower bound for 1-spanner of the uniform line metric $L_{n}=\{1,2,\ldots,n\}$ with hop-diameter $k$. Due to the inductive nature of our argument, we prove a stronger version of the statement, which considers 1-spanners that could potentially use points outside of the given line metric.
@@
We do so by arguing that any 1-spanner of $L_{n}$ has a large minor. It is well known that a graph with $K_{t}$ as a minor has treewidth at least t1t-1.
@@ 3.2.1 Hop-diameter 2
For every $n\geq 1$ and every $n$-vertex line metric $L_{n}$, every 1-spanner with hop-diameter 2 has ${K}_{\lfloor\log{n}\rfloor+1}$ as a minor.
@@
Let $H$ be a 1-spanner for $L_{n}$ with hop-diameter 2. Split $L_{n}$ into two consecutive parts, $L_{1}$ and $L_{2}$, of sizes $\lfloor n/2\rfloor$ and $\lceil n/2\rceil$, respectively. Let $H_{1}$ and $H_{2}$ be subgraphs of $H$ induced on $L_{1}$ and $L_{2}$, respectively. From the induction hypothesis, $H_{1}$ and $H_{2}$ have $K_{\log{\lfloor n/2\rfloor}}$ and $K_{\log{\lceil n/2\rceil}}$ as minors, respectively.
@@ 3.2.2 Hop-diameter $k\geq 4$
For every $n\geq 1$, every even $4\leq k\leq\frac{2}{\ln(2e)}\ln\log{n}$, and every $n$-vertex line metric $L$, every 1-spanner with hop-diameter $k$ has treewidth at least $c_{1}k\log^{2/k}n-1$, where $c_{1}$ is an absolute constant.
@@
Split $L$ into consecutive sets of points $L_{1},L_{2},\ldots,L_{\ell_{k}}$ of size $\lfloor n/\ell_{k}\rfloor$ each and ignore the remaining points. Let $H$ be a 1-spanner with hop-diameter $k$ for $L_{n}$. Our goal is to show that the size of a clique minor of $H$ can be lower bounded by the following recurrence.
@@
For every $L_{i}$ and $L_{j}$ the path between a non-global point in $L_{i}$ and a non-global point in $L_{j}$ must have the first (resp., last) edge inside $L_{i}$ (resp., $L_{j}$). Let $H^{\prime}$ be obtained from $H$ by contracting each $L_{i}$ into a single vertex. (Clearly, $H[L_{i}]$ is connected.) Let $L^{\prime}$ be the line metric obtained from $L_{n}$ by contracting every $L_{i}$ into a single point. Then $H^{\prime}$ is a (k2)(k-2)-hop spanner of $L^{\prime}$ with stretch 1. Thus, $H^{\prime}$ has a minor of size $W_{k-2}(\ell_{k})\geq c_{1}(k-2)\log^{2/(k-2)}\ell_{k}-1=c_{1}k\log^{2/k}{n}-1$.
@@
Let $\{L_{l},L_{r}\}=L\setminus L_{i}$, so that $L_{l}$ (resp., $L_{r}$) is on the left (resp., right) of $L_{i}$. (Possibly $L_{l}=\emptyset$ or $L_{r}=\emptyset$.) At least $|L_{i}|/2$ points in $L_{i}$ have edges to either $L_{l}$ or $L_{r}$. Without loss of generality, we assume the former. Let $L^{\prime}_{i}\subseteq L_{i}$ be the subset of points that have edges to $L_{l}$ and let $H_{i}$ be the subgraph of $H$ restricted to preserving distances in $L_{i}$. Inductively, $H_{i}$ has a clique minor of size at least $W_{k}(n/(2\ell_{k}))$. (Since $H_{i}$ is a $1$-spanner, it does not include any point outside of $L_{i}$.) Then ${H}_{i}$ and ${L}_{l}$ are vertex-disjoint (because we are considering $1$-spanners) and hence their union has a clique minor of size $W_{k}(n/(2\ell_{k}))+1$. Thus, $W_{k}(n)$ satisfies eq.˜5, which we lower bound next.
@@
For every $n\geq 1$, every $k>\frac{2}{\ln(2e)}\ln\log{n}$, every 1-spanner with hop-diameter $k$ for an $n$-vertex line metric has treewidth at least $c_{2}(\log\log{n})^{2}/k$, for an absolute constant $c_{2}$.
@@
4 Bounded arboricity tree covers

### v2

@@ 2 Preliminaries
A tree decomposition of $G=(V,E)$ is a tree $T$ where each node is associated with a subset of $V$ called a bag. The bags $X_{1},\ldots,X_{\ell}$ satisfy: (i) $\cup_{i=1}^{\ell}X_{i}=V$, (ii) for every vertex $v\in V$, the collection of bags containing it forms a connected subtree of $T$, and (iii) for every edge $(u,v)\in E$, there is a bag containing it. The width of $T$ is $\max_{i=1}^{\ell}|X_{i}|-1$; the treewidth of $G$ is the minimum width among all possible tree decompositions of $G$. If a graph $G$ contains $K_{h}$ as a minor, then the treewidth of $G$ is at least h1h-1.
@@
The arboricity of $G=(V,E)$ is defined as $\max\lceil\frac{|E(H)|}{|V(H)|-1}\rceil$, where the maximum is taken over all subgraphs $H$ of $G$ with at least two vertices.
Spanners.
Given a metric space $M_{X}=(X,\delta)$, which we view as a complete graph $(X,\binom{X}{2})$, a $t$-spanner of $M_{X}$ is a subgraph $H=(X,E)$ with $E\subseteq\binom{X}{2}$, such that for every two vertices $x,y\in X$, their distance in $H$ is at most $t\cdot\delta(x,y)$. The path $P_{x,y}$ realizing this distance is called a $t$-spanner path. The hop-diameter of a spanner is the minimum $k$ such that between every two vertices $x,y\in X$ there is a $t$-spanner path with at most $k$ edges.
Tree covers.
Given a metric space $M_{X}=(X,\delta)$, a tree cover of $M_{X}$ with stretch $t$ is a collection of trees such that for every two vertices $x,y\in X$ we have $\delta(x,y)\leq\min_{T\in\mathcal{T}}\delta_{T}(x,y)\leq t\cdot\delta(x,y)$, where $\delta_{T}(x,y)$ is the distance between $x$ and $y$ in $T$.
@@
For all $k\geq 0$, the Ackermann functions $A(k,n)$ and $B(k,n)$ are defined as follows:
@@
For all $k\geq 0$ and $n\geq 0$, the inverse Ackermann function $\alpha_{k}(n)$ is defined as follows:
$\alpha_{2k}(n):=\min\{s\geq 0:A(k,s)\geq n\}$
$\alpha_{2k+1}(n):=\min\{s\geq 0:B(k,s)\geq n\}$
@@
Compact routing schemes.
A compact routing scheme is a distributed algorithm that can deliver packets between any source node and any destination node in the network. Each node in the network has its own routing table, which stores local routing information, as well as a unique label. During a preprocessing phase, the network is initialized so that each node is assigned a routing table and a label. In the labeled model, labels are chosen by the designer (typically of size $\mathrm{poly}(\log n)$), while in the name-independent model, labels are chosen adversarially.
3 Low treewidth shortcuttings
In this section, we show tight tradeoff between treewidth and hop-diameter for tree shortcuttings. In particular, the upper bound (Section˜1.1) is proved in Section˜3.1 and the matching lower bound (Section˜1.1) is proved in Section˜3.2.
@@ 3.1 Upper bound
In this section we prove Section˜1.1.
@@ 3.1.1 Hop-diameter 2
For every tree $T$ there is a shortcutting $H_{2}$ with hop-diameter 2 and treewidth $O(\log{n})$.
@@
Let $\mathcal{T}_{1},\ldots,\mathcal{T}_{g}$ be the tree decompositions of shortcuttings constructed for the subtrees in $T\setminus v$. Create a new bag $B$ containing only the vertex $v$ and add $v$ to every bag in each $\mathcal{T}_{1},\ldots,\mathcal{T}_{g}$. The tree decomposition of $H_{2}$ is obtained by connecting $B$ to the roots of each of $\mathcal{T}_{1},\ldots,\mathcal{T}_{g}$. The treewidth satisfies recurrence $W_{2}(1)=0$ and $W_{2}(n)=W_{2}(n/2)+1$, which has solution $W_{2}(n)=O(\log{n})$.
@@ 3.1.2 Hop-diameter 3
For every tree $T$ there is a shortcutting $H_{3}$ with hop-diameter 3 and treewidth $O(\log{n}/\log\log{n})$.
@@
Let $\mathcal{T}_{1},\ldots,\mathcal{T}_{g}$ be the tree decompositions of shortcuttings constructed for the trees in $T\setminus X$. For every $\mathcal{T}_{i}$, let $u$ and $v$ be the two vertices from $X$ adjacent to the corresponding subtree $W_{i}$ in $T\setminus X$. (Lemma˜3.1 guarantees that there is at most two such vertices; it is possible that $u=v$.) Add $u$ and $v$ to each bag in $\mathcal{T}_{i}$. Construct a new bag $B$ containing all the vertices in $X$ and connect it to the roots of each $\mathcal{T}_{1},\ldots,\mathcal{T}_{g}$. The treewidth of $H_{3}$ satisfies $W_{3}(n)\leq n$ for $n\leq\ell_{3}$ and $W_{3}(n)\leq W_{3}(n/\ell_{3})+2$ for $n>\ell_{3}$. Recall that $\ell_{3}$ is fixed and does not change across different levels of recursion. The recurrence satisfies $W_{3}(n)=O(\log{n}/\log\log{n})$.
@@ 3.1.3 Hop-diameter $k\geq 4$
For every tree $T$ and every $k\geq 4$ there is a shortcutting $H_{k}$ with hop-diameter $k$ and treewidth $O(k\log^{2/k}{n})$ for even values of $k$ and $O(k(\frac{\log{n}}{\log\log{n}})^{2/(k-1)})$ for odd values of $k$.
@@
Next, we analyze the treewidth of $H_{k}$. Let $\mathcal{T}_{1},\ldots,\mathcal{T}_{g}$ be the tree decompositions of shortcuttings constructed for the trees in $T\setminus X$. For every $\mathcal{T}_{i}$, let $u$ and $v$ be the two vertices from $X$ adjacent to the corresponding subtree $W_{i}$ in $T\setminus X$. (Lemma˜3.1 guarantees that there is at most two such vertices; it is possible that $u=v$.) Add $u$ and $v$ to each bag in $\mathcal{T}_{i}$. Let $\mathcal{T}_{X}$ be the tree decomposition of $T_{X}$, where $T_{X}$ is defined as in the previous paragraph. Since $\mathcal{T}_{X}$ is a valid tree decomposition and $T_{X}$ contains an edge $(u,v)$, then $\mathcal{T}_{X}$ contains a bag $L_{u,v}$ with both $u$ and $v$. Connect the root of $\mathcal{T}^{\prime}$ to $L_{u,v}$. This concludes the description of the tree decomposition $\mathcal{T}$ of $T$. The treewidth of $\mathcal{T}$ satisfies the recurrence $W_{k}(n)=W_{k-2}(n)$ for $n\leq\ell_{k}$ and $W_{k}(n)\leq\max(W_{k-2}(\ell_{k}),W_{k}(n/\ell_{k})+2)$ otherwise.
@@
To argue that $H_{k}$ is a shortcutting with hop-diameter $k$, consider the last step of the construction where both $u$ and $v$ were in the same tree $T$. If $u$ and $v$ were considered in the same base case, then $T$ is equipped with a recursive construction with parameter k2k-2. By induction, there is a shortcutting path between them with hop-diameter k2k-2. Otherwise, let $W_{u}$ (resp., $W_{v}$) be the subtree in $T\setminus X$ and let $u^{\prime}$ (resp., $v^{\prime}$) be the vertex in $X$ that is incident on $W_{u}$ (resp., $W_{v}$). By construction, there is a shortcutting path between $u^{\prime}$ and $v^{\prime}$ with at most k2k-2 hops. (The cases where $u^{\prime}=u$ or $v^{\prime}=v$ are handled similarly.) ∎
@@ 3.2 Lower bound
We show the treewidth lower bound for shortcutting of the path $L_{n}=\{1,2,\ldots,n\}$ with hop-diameter $k$. Due to the inductive nature of our argument, we prove a stronger version of the statement, which considers shortcuttings that could potentially use points outside of the given path.
@@
We do so by arguing that any shortcutting of $L_{n}$ has a large minor. It is well known that a graph with $K_{t}$ as a minor has treewidth at least t1t-1.
@@ 3.2.1 Hop-diameter 2
For every $n\geq 1$ and every $n$-vertex path $L_{n}$, every shortcutting with hop-diameter 2 has ${K}_{\lfloor\log{n}\rfloor+1}$ as a minor.
@@
Let $H$ be a shortcutting for $L_{n}$ with hop-diameter 2. Split $L_{n}$ into two consecutive parts, $L_{1}$ and $L_{2}$, of sizes $\lfloor n/2\rfloor$ and $\lceil n/2\rceil$, respectively. Let $H_{1}$ and $H_{2}$ be subgraphs of $H$ induced on $L_{1}$ and $L_{2}$, respectively. From the induction hypothesis, $H_{1}$ and $H_{2}$ have $K_{\log{\lfloor n/2\rfloor}}$ and $K_{\log{\lceil n/2\rceil}}$ as minors, respectively.
@@ 3.2.2 Hop-diameter $k\geq 4$
For every $n\geq 1$, every even $4\leq k\leq\frac{2}{\ln(2e)}\ln\log{n}$, and every $n$-vertex path $L$, every shortcutting with hop-diameter $k$ has treewidth at least $c_{1}k\log^{2/k}n-1$, where $c_{1}$ is an absolute constant.
@@
Split $L$ into consecutive sets of points $L_{1},L_{2},\ldots,L_{\ell_{k}}$ of size $\lfloor n/\ell_{k}\rfloor$ each and ignore the remaining points. Let $H$ be a shortcutting with hop-diameter $k$ for $L_{n}$. Our goal is to show that the size of a clique minor of $H$ can be lower bounded by the following recurrence.
@@
For every $L_{i}$ and $L_{j}$ the path between a non-global point in $L_{i}$ and a non-global point in $L_{j}$ must have the first (resp., last) edge inside $L_{i}$ (resp., $L_{j}$). Let $H^{\prime}$ be obtained from $H$ by contracting each $L_{i}$ into a single vertex. (Clearly, $H[L_{i}]$ is connected.) Let $L^{\prime}$ be the path obtained from $L_{n}$ by contracting every $L_{i}$ into a single point. Then $H^{\prime}$ is a (k2)(k-2)-hop shortcutting of $L^{\prime}$ with stretch 1. Thus, $H^{\prime}$ has a minor of size $W_{k-2}(\ell_{k})\geq c_{1}(k-2)\log^{2/(k-2)}\ell_{k}-1=c_{1}k\log^{2/k}{n}-1$.
Figure 1: An illustration of Case 1. Every set $L_{i}$ contains a non-global point. Any $k$-hop path between two non-global points must have the first and the last edge inside the corresponding sets. By contracting every $L_{i}$ into a single point, we obtain a new path $L^{\prime}$. Any shortcutting with hop-diameter $k$ for $L$ must have hop-diameter k2k-2 for $L^{\prime}$. This allows us to use the induction hypothesis for k2k-2. In the illustration above the blue path consisting of 4 hops in $L$ becomes a 2-hop path in $L^{\prime}$.
@@
Let $\{L_{l},L_{r}\}=L\setminus L_{i}$, so that $L_{l}$ (resp., $L_{r}$) is on the left (resp., right) of $L_{i}$. (Possibly $L_{l}=\emptyset$ or $L_{r}=\emptyset$.) At least $|L_{i}|/2$ points in $L_{i}$ have edges to either $L_{l}$ or $L_{r}$. Without loss of generality, we assume the former. Let $L^{\prime}_{i}\subseteq L_{i}$ be the subset of points that have edges to $L_{l}$ and let $H_{i}$ be the subgraph of $H$ restricted to preserving distances in $L_{i}$. Inductively, $H_{i}$ has a clique minor of size at least $W_{k}(n/(2\ell_{k}))$. (Since $H_{i}$ is a shortcutting, it does not include any point outside of $L_{i}$.) Then ${H}_{i}$ and ${L}_{l}$ are vertex-disjoint (because we are considering shortcuttings) and hence their union has a clique minor of size $W_{k}(n/(2\ell_{k}))+1$. Thus, $W_{k}(n)$ satisfies eq.˜5, which we lower bound next.
Figure 2: An illustration of Case 2. $L_{2}$ is global, meaning that every vertex is adjacent to an edge going outside of $L_{2}$. At least half of these edges are adjacent on the vertices in $L_{r}$. By contracting $L_{r}$ and using the clique minor of $L_{2}$, we obtain the desired clique minor.
@@
For every $n\geq 1$, every $k>\frac{2}{\ln(2e)}\ln\log{n}$, every 1-shortcutting with hop-diameter $k$ for an $n$-vertex path has treewidth at least $c_{2}(\log\log{n})^{2}/k$, for an absolute constant $c_{2}$.
@@
4 Low arboricity shortcuttings